# The Complexity of Sequential Prediction in Dynamical Systems

### v1

@@ G.1 Discrete Linear Systems
(of Theorem 3.6) We first prove the lowerbound. It suffices to show that
$\inf_{\text{Deterministic }\mathcal{A}}\,\,\texttt{M}_{{\mathcal{A}}}(T,\mathcal{F})\geq r+1.$
@@
Let $\mathcal{A}$ be any deterministic algorithm and $\{e_{0},e_{1},\ldots,e_{r}\}$ be arbitrary $r+1$ standard basis on $\mathbb{R}^{n}$. This set exists because $n\geq r+1$. Consider a stream such that $x_{0}=e_{0}$, $x_{t}=\{-e_{t},e_{t}\}\backslash\mathcal{A}(x_{<t})$ for all $t\in[r]$, and $x_{r+1}=\{-e_{1},e_{1}\}\backslash\mathcal{A}(x_{<r+1}).$ Here, $\forall t\in[r]$, we choose $x_{t}$ to be an element of the set $\{-e_{t},e_{t}\}$ other than $\mathcal{A}$’s prediction on round $t$. Similarly, $x_{r+1}$ is chosen among $\{-e_{1},e_{1}\}$. We can define a stream using $\mathcal{A}$ because such a stream can be simulated before the game starts as $\mathcal{A}$ is deterministic. Let $\{\sigma_{t}\}_{t=0}^{r+1}\in\{-1,1\}^{r+2}$ such that $(x_{0},\ldots,x_{r+1})=(\sigma_{0}\,e_{0},\sigma_{1}e_{1},\sigma_{2}e_{2},\ldots,\sigma_{r}e_{r},\sigma_{r+1}e_{1})$. Consider the integer-valued matrix $W^{\star}=\sum_{t=1}^{r}\sigma_{t-1}\sigma_{t}\,e_{t}\otimes e_{t-1}+\sigma_{r}\sigma_{r+1}\,e_{1}\otimes e_{r}$. For $r+2\leq t\leq T$, one can define the stream to be $x_{t}=W^{\star}x_{t-1}$. Note that the $\text{image}(W^{\star})\subseteq\text{span}(\{e_{1},\ldots,e_{r}\})$. Thus, $\text{rank}(W^{\star})\leq r$ and $W^{\star}\in\mathcal{F}$. Finally, since $W^{\star}$ satisfies $W^{\star}(\sigma_{t-1}e_{t-1})=\sigma_{t}\,e_{t}$ for $t\in[r]$, $W^{\star}(\sigma_{r}e_{r})=\sigma_{r+1}e_{1}$, and $W^{\star}x_{t-1}=x_{t}$ for $t\geq r+2$, the stream is realizable. By the definition of the stream, $\mathcal{A}$ makes mistakes on all rounds $t\in\{1,\ldots,r+1\}$. Thus, we have $\texttt{M}_{{\mathcal{A}}}(T,\mathcal{F})\geq r+1$.
@@
To prove the upperbound, fix $T\in\mathbbm{N}$ and suppose that $T\geq r+1$ (otherwise $\texttt{C}_{T}(\mathcal{F})\leq T\leq r+1$). Consider a trajectory tree $\mathcal{T}$ of depth $T$ shattered by $\mathcal{F}$. We first show that, for any path $\sigma\in\{-1,1\}^{T}$ down $\mathcal{T}$ and $t\in[T-1]$, if $\mathcal{T}_{t+1}((\sigma_{\leq t},-1))\neq\mathcal{T}_{t+1}((\sigma_{\leq t},+1))$, then $\mathcal{T}_{t}(\sigma_{\leq t})\notin\text{span}(\{\mathcal{T}_{0},\mathcal{T}_{1}(\sigma_{\leq 1}),...,\mathcal{T}_{t-1}(\sigma_{\leq t-1}\})$. To prove the contraposition of this statement, suppose $\mathcal{T}_{t}(\sigma_{\leq t})\in\text{span}(\{\mathcal{T}_{0},\mathcal{T}_{1}(\sigma_{\leq 1}),...,\mathcal{T}_{t-1}(\sigma_{\leq t-1}\})$. Then, there exists constants $a_{0},...,a_{t-1}\in\mathbbm{Z}$ such that $a_{0}\mathcal{T}_{0}+...,+a_{t-1}\mathcal{T}_{t-1}(\sigma_{\leq t-1})=\mathcal{T}_{t}(\sigma_{\leq t})$. Thus, every matrix $W$ consistent with the sequence $\mathcal{T}_{0},\mathcal{T}_{1}(\sigma_{\leq 1}),...,\mathcal{T}_{t-1}(\sigma_{\leq t-1})$ satisfies $W\mathcal{T}_{t}(\sigma_{\leq t})=W(a_{0}\mathcal{T}_{0}+...,+a_{t-1}\mathcal{T}_{t-1}(\sigma_{\leq t-1}))=a_{0}\mathcal{T}_{1}(\sigma_{\leq 1})+...,+a_{t-1}\mathcal{T}_{t}(\sigma_{\leq t})$. This implies that the path $\sigma_{\leq t+1}$ is shattered only if $\mathcal{T}_{t+1}((\sigma_{\leq t},-1))=\mathcal{T}_{t+1}((\sigma_{\leq t},+1))$.
@@
Next, we show that the branching factor of every path in $\mathcal{T}$ can be at most $r+1$. Suppose, for the sake of contradiction, there exists a path $\sigma\in\{-1,1\}^{T}$ whose branching factor is $k>r+1$. We are guaranteed an increasing sequence of time points $t_{1},t_{2},\ldots,t_{k}$ in $\{0,\ldots,T-1\}$ such that $\mathcal{T}_{t_{i}+1}((\sigma_{\leq t_{i}},-1))\neq\mathcal{T}_{t_{i}+1}((\sigma_{\leq t_{i}},+1))$ for all $i\in[k]$. Since $t_{2}\geq 1$, by our argument above, we are guaranteed that $\mathcal{T}_{t_{i}}(\sigma_{\leq t_{i}})\notin\text{span}(\{\mathcal{T}_{0},\ldots,\mathcal{T}_{t_{i}-1}(\sigma_{\leq t_{i}-1})\})$ for all $i\in\{2,\ldots,k\}$. This further implies that the set $S:=\{\mathcal{T}_{t_{2}}(\sigma_{\leq t_{2}}),\ldots,\mathcal{T}_{t_{k}}(\sigma_{\leq t_{k}})\}$ is a linearly independent set. Note that $|S|=k-1>r$. Let $W_{\sigma}$ be the matrix such that the associated function $f_{\sigma}\in\mathcal{F}$ shatters this path $\sigma$. By the definition of the tree, we must have that $S\subseteq\text{image}(W_{\sigma})$. Since $\text{rank}(W_{\sigma})\leq r$, the set $\text{image}(W_{\sigma})$ must be a subset of a vectorspace of dimension $\leq r$. However, this contradicts the fact that $S$ contains at least $r+1$ linearly independent vectors in $\mathbb{R}^{n}$. Thus, the branching factor of every path in $\mathcal{T}$ is at most $r+1$ and $\texttt{B}(\mathcal{T})\leq r+1$. Since $\mathcal{T}$ is arbitrary, we have $\texttt{C}_{T}(\mathcal{F})\leq r+1$. ∎
G.2 Linear Boolean Networks
Proof.
@@ G.2 Linear Boolean Networks
(of (i) in Theorem 3.7) We claim that it is without loss of generality to consider the function class $\mathcal{G}=\{x\mapsto Mx\,(\text{mod }2)\,:\,M\in\{0,1\}^{n\times n}\}\subset\mathcal{F}$. Indeed, for every $f\in\mathcal{F}$, there exists a $g\in\mathcal{G}$ such that $f(x)=g(x)$ for all $x\in\mathcal{X}$. To see this, let $W\in\mathbb{Z}^{n\times n}$ be such that $f(x)=Wx(\text{mod }2)$ and fix some $x\in\{0,1\}^{n}$. Then, observe that $f(x)=Wx(\text{mod }2)=(W(\text{mod }2)x)(\text{mod }2)=Mx(\text{mod }2)$ where $M=W(\text{mod }2)\in\{0,1\}^{n\times n}$. Thus the function $g\in\mathcal{G}$ parameterized by $M$ matches $f$ everywhere on $\mathcal{X}$. We now prove that $\texttt{C}_{T}(\mathcal{G})=n$. In the lowerbound, we will use the fact that for a system of $n$ linearly independent equations with $n$ free variables defined over the integers modulo $2$, there exists a unique solution.
@@
Starting with the upperbound, fix $T\in\mathbbm{N}$ and suppose that $T\geq n$ (as otherwise $\texttt{C}_{T}(\mathcal{G})\leq T\leq n$). Consider a trajectory tree $\mathcal{T}$ of depth $T$ shattered by $\mathcal{G}$. Let $\sigma\in\{-1,1\}^{T}$ denote an arbitrary path down $\mathcal{T}$. We first claim that for $t\geq 1$ if $\mathcal{T}_{t}(\sigma_{\leq t})$ is linearly dependent modulo $2$ on the preceding sequence of states $\mathcal{T}_{0},\mathcal{T}_{1}(\sigma_{\leq 1}),...,\mathcal{T}_{t-1}(\sigma_{<t})$, then $\mathcal{T}_{t+1}((\sigma_{\leq t},-1))=\mathcal{T}_{t+1}((\sigma_{\leq t},+1)).$ To see this, suppose that $\mathcal{T}_{t}(\sigma_{\leq t})$ is linearly dependent modulo $2$ on the preceding sequence of states $\mathcal{T}_{0},\mathcal{T}_{1}(\sigma_{\leq 1}),...,\mathcal{T}_{t-1}(\sigma_{<t})$. Then, there exists constants $a_{0},...,a_{t-1}\in\{0,1\}$ such that $a_{0}\mathcal{T}_{0}+...+a_{t-1}\mathcal{T}_{t-1}(\sigma_{<t})(\text{mod }2)=\mathcal{T}_{t}(\sigma_{\leq t}).$ Thus, every function $g\in\mathcal{G}$ consistent with the sequence $\mathcal{T}_{0},\mathcal{T}_{1}(\sigma_{\leq 1}),...,\mathcal{T}_{t-1}(\sigma_{<t})$ outputs $a_{0}\mathcal{T}_{1}(\sigma_{\leq 1})+...+a_{t-1}\mathcal{T}_{t}(\sigma_{\leq t})(\text{mod }2)$ on $\mathcal{T}_{t}(\sigma_{\leq t})$, implying that the paths $(\sigma_{\leq t},-1)$ and $(\sigma_{\leq t},+1)$ are shattered only if $\mathcal{T}_{t+1}((\sigma_{\leq t},-1))=\mathcal{T}_{t+1}((\sigma_{\leq t},+1))=a_{0}\mathcal{T}_{1}(\sigma_{\leq 1})+...+a_{t-1}\mathcal{T}_{t}(\sigma_{\leq t})(\text{mod }2).$ As a consequence, for $t\geq 1$, if $\mathcal{T}_{t+1}((\sigma_{\leq t},-1))\neq\mathcal{T}_{t+1}((\sigma_{\leq t},+1))$, then $\mathcal{T}_{t}(\sigma_{\leq t})$ is linearly independent modulo $2$ of its preceding states. Next, we claim that there can be at most $n-1$ timepoints $t_{1},...,t_{n-1}\in[T-1]$ such that $\mathcal{T}_{t_{i}}(\sigma_{\leq t_{i}})$ is linearly independent of its preceeding sequence of states $\mathcal{T}_{0},\mathcal{T}_{1}(\sigma_{\leq 1}),...,\mathcal{T}_{t_{i}-1}(\sigma_{<t_{i}})$. Indeed, suppose for sake of contradiction there exists $n$ timepoints $t_{1},...,t_{n}\in[T-1]$ such that $\mathcal{T}_{t_{i}}(\sigma_{\leq t_{i}})$ is linearly independent of its preceeding states. Then, note that the following set of $n+1$ states $\{\mathcal{T}_{0},\mathcal{T}_{t_{1}}(\sigma_{\leq t_{1}}),...,\mathcal{T}_{t_{n}}(\sigma_{\leq t_{n}})\}$ are linearly independent modulo $2$. This is a contradiction since $\mathcal{X}$ is $n$-dimensional. Combining the two claims, we get that
@@
Since $\sigma\in\{-1,1\}^{T}$ is arbitrary, we get that $\texttt{B}(\mathcal{T})\leq n$. Finally, since $\mathcal{T}$ is arbitrary, we have that $\texttt{C}_{T}(\mathcal{F})\leq n$ which completes the proof of the upperbound.
To prove the lowerbound, let $\mathcal{A}$ be any deterministic algorithm. Let =$\mathcal{E}=\{e_{0},...,e_{n-1}\}\subset\mathcal{X}$ be the standard basis over $\mathbb{R}^{n}$. Consider the stream where we pick $x_{0}=e_{0}$, for all $t\in[n-2]$, we pick $x_{t}\in\mathcal{E}\setminus\{x_{0},...,x_{t-1},\mathcal{A}(x_{<t})\}$. Let $e=\mathcal{E}\setminus\{x_{0},...,x_{n-2}\}$ be the remaining basis. Pick $x_{n-1}\in\{e,e+e_{0}\}\setminus\{\mathcal{A}(x_{<n-1})\}$. Finally, pick $x_{n}\neq\mathcal{A}(x_{<n})$. We can define a stream using $\mathcal{A}$ because it is deterministic and thus can be simulated before the game begins. Next, we show that $x_{0},...,x_{n}$ is a realizable stream. This follows from the fact that $x_{0},...,x_{n-1}$ are linearly independent by definition, and thus there exists a function $g\in\mathcal{G}$ such that $g(x_{t-1})=x_{t}$ for all $t\in[n]$. Moreover, by definition of the stream, $\mathcal{A}$ makes a mistake in every round. Thus, $\texttt{M}_{\mathcal{A}}(T,\mathcal{F})\geq n$. Using the fact that $\texttt{M}_{\mathcal{A}}(T,\mathcal{F})\leq\texttt{C}_{T}(\mathcal{F})$ completes the proof. ∎
Proof.
(of (ii) in Theorem 3.7) Starting with the lowerbound, it suffices to show that
$\inf_{\text{Deterministic }\mathcal{A}}\texttt{M}_{{\mathcal{A}}}(T,\mathcal{F})\geq n.$
@@
So, we obtain $\mathbbm{1}\{W^{\star}x_{t-1}>0\}=x_{0}-(e_{i_{1}}+\ldots+e_{i_{t}})>0=x_{t}$. Finally, since $x_{n-1}=e_{i_{n}}$ and $W^{\star}e_{i_{n}}=x_{n}$, we have $\mathbbm{1}\{W^{\star}x_{n-1}>0\}=x_{n}$. By the definition of the stream, $\mathcal{A}$ makes mistakes in every round on this stream. Thus, $\texttt{M}_{{\mathcal{A}}}(T,\mathcal{F})\geq n$.
As for the upperbound, since |$|\{0,1\}^{n\times n}|=2^{n^{2}}$, we have $\texttt{L}(\mathcal{F})\leq\log(2^{n^{2}})=n^{2}$. Thus, by Theorem 3.5 (iii), we must have $\texttt{C}_{T}(\mathcal{F})\leq n^{2}$. ∎
Appendix H Proof of Theorem 4.1
Proof.
@@
where expectation is taken with respect to the randomness of $\mathcal{B}$ and $R$ is the expected regret of the $\mathcal{B}$. This shows that $\texttt{R}^{\texttt{MK}}_{\mathcal{A}}(T,\mathcal{F})\leq R$. Taking $\mathcal{B}$ to be $\mathbb{A}_{\text{AG}}$ defined in [Hanneke et al., 2023, Section 3], Theorem 4 in [Hanneke et al., 2023] implies that $R\leq\texttt{L}(\mathcal{F})+\sqrt{T\,\texttt{L}(\mathcal{F})\log T}$, matching the claimed upperbound in Theorem 4.1. ∎
Proof.
(of lowerbound of $\frac{\texttt{L}(\mathcal{F})}{18}$ in Theorem 4.1) Define $d:=\texttt{L}(\mathcal{F})$ and let $\mathcal{T}$ be the Littlestone tree of depth $d$ shattered by $\mathcal{F}$. Let $(\mathcal{T}_{0},\ldots,\mathcal{T}_{d-1})$ be the sequence of node-labeling functions and $(Y_{1},\ldots,Y_{d})$ be the sequence of edge labeling functions of the shattered tree $\mathcal{T}$. To construct a stream, take $r=\lfloor\frac{d+2}{3}\rfloor$ and define a random sequence $\{n_{i}\}_{i=1}^{r}$ such that
@@
$\texttt{R}^{\texttt{MK}}_{\mathcal{A}}(T,\mathcal{F})\geq\frac{7r-4}{6}-(r-1)=\frac{7r-4-6r+6}{6}=\frac{r+2}{6}\geq\frac{(d+2)/3-1+2}{6}\geq\frac{d}{18}.$
This completes our proof.∎
Proof.
@@
$\begin{split}\texttt{R}^{\texttt{MK}}_{\mathcal{A}}(T,\mathcal{F})&=\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{A}(x_{<t})\neq x_{t}\}\right]-\mathbb{E}\left[\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\mathbbm{1}\{f(x_{t-1})\neq x_{t}\}\right]\geq\frac{1}{2}\mathbb{E}\left[\left|\sum_{t=1}^{T}\sigma_{t}\mathbbm{1}\{x_{t-1}\notin S\}\right|\right].\end{split}$
We now lowerbound the Rademacher sum above by closely following the proof of Khinchine’s inequality [Cesa-Bianchi and Lugosi, 2006, Lemma 8.1].

### v2

@@ Appendix E Proofs for Linear Systems
(of (i) in Theorem 3.6) We first prove the lower bound. It suffices to show that
$\inf_{\text{Deterministic }\mathcal{A}}\,\,\operatorname{M}_{{\mathcal{A}}}(T,\mathcal{F})\geq r+1.$
@@
Let $\mathcal{A}$ be any deterministic algorithm and $\{e_{0},e_{1},\ldots,e_{r}\}$ be arbitrary $r+1$ standard basis on $\mathbb{R}^{n}$. This set exists because $n\geq r+1$. Consider a stream such that $x_{0}=e_{0}$, $x_{t}=\{-e_{t},e_{t}\}\backslash\mathcal{A}(x_{<t})$ for all $t\in[r]$, and $x_{r+1}=\{-e_{1},e_{1}\}\backslash\mathcal{A}(x_{<r+1}).$ Here, $\forall t\in[r]$, we choose $x_{t}$ to be an element of the set $\{-e_{t},e_{t}\}$ other than $\mathcal{A}$’s prediction on round $t$. Similarly, $x_{r+1}$ is chosen among $\{-e_{1},e_{1}\}$. We can define a stream using $\mathcal{A}$ because such a stream can be simulated before the game starts as $\mathcal{A}$ is deterministic. Let $\{\sigma_{t}\}_{t=0}^{r+1}\in\{-1,1\}^{r+2}$ such that $(x_{0},\ldots,x_{r+1})=(\sigma_{0}\,e_{0},\sigma_{1}e_{1},\sigma_{2}e_{2},\ldots,\sigma_{r}e_{r},\sigma_{r+1}e_{1})$. Consider the integer-valued matrix $W^{\star}=\sum_{t=1}^{r}\sigma_{t-1}\sigma_{t}\,e_{t}\otimes e_{t-1}+\sigma_{r}\sigma_{r+1}\,e_{1}\otimes e_{r}$. For $r+2\leq t\leq T$, one can define the stream to be $x_{t}=W^{\star}x_{t-1}$. Note that the $\text{image}(W^{\star})\subseteq\text{span}(\{e_{1},\ldots,e_{r}\})$. Thus, $\text{rank}(W^{\star})\leq r$ and $W^{\star}\in\mathcal{F}$. Finally, since $W^{\star}$ satisfies $W^{\star}(\sigma_{t-1}e_{t-1})=\sigma_{t}\,e_{t}$ for $t\in[r]$, $W^{\star}(\sigma_{r}e_{r})=\sigma_{r+1}e_{1}$, and $W^{\star}x_{t-1}=x_{t}$ for $t\geq r+2$, the stream is realizable. By the definition of the stream, $\mathcal{A}$ makes mistakes on all rounds $t\in\{1,\ldots,r+1\}$. Thus, we have $\operatorname{M}_{{\mathcal{A}}}(T,\mathcal{F})\geq r+1$.
@@
To prove the upper bound, fix $T\in\mathbbm{N}$ and suppose that $T\geq r+1$ (otherwise $\operatorname{C}_{T}(\mathcal{F})\leq T\leq r+1$). Consider a trajectory tree $\mathcal{T}$ of depth $T$ shattered by $\mathcal{F}$. We first show that, for any path $\sigma\in\{-1,1\}^{T}$ down $\mathcal{T}$ and $t\in[T-1]$, if $\mathcal{T}_{t+1}((\sigma_{\leq t},-1))\neq\mathcal{T}_{t+1}((\sigma_{\leq t},+1))$, then $\mathcal{T}_{t}(\sigma_{\leq t})\notin\text{span}(\{\mathcal{T}_{0},\mathcal{T}_{1}(\sigma_{\leq 1}),...,\mathcal{T}_{t-1}(\sigma_{\leq t-1}\})$. To prove the contraposition of this statement, suppose $\mathcal{T}_{t}(\sigma_{\leq t})\in\text{span}(\{\mathcal{T}_{0},\mathcal{T}_{1}(\sigma_{\leq 1}),...,\mathcal{T}_{t-1}(\sigma_{\leq t-1}\})$. Then, there exists constants $a_{0},...,a_{t-1}\in\mathbbm{Z}$ such that $a_{0}\mathcal{T}_{0}+...,+a_{t-1}\mathcal{T}_{t-1}(\sigma_{\leq t-1})=\mathcal{T}_{t}(\sigma_{\leq t})$. Thus, every matrix $W$ consistent with the sequence $\mathcal{T}_{0},\mathcal{T}_{1}(\sigma_{\leq 1}),...,\mathcal{T}_{t-1}(\sigma_{\leq t-1})$ satisfies $W\mathcal{T}_{t}(\sigma_{\leq t})=W(a_{0}\mathcal{T}_{0}+...,+a_{t-1}\mathcal{T}_{t-1}(\sigma_{\leq t-1}))=a_{0}\mathcal{T}_{1}(\sigma_{\leq 1})+...,+a_{t-1}\mathcal{T}_{t}(\sigma_{\leq t})$. This implies that the path $\sigma_{\leq t+1}$ is shattered only if $\mathcal{T}_{t+1}((\sigma_{\leq t},-1))=\mathcal{T}_{t+1}((\sigma_{\leq t},+1))$.
@@
Next, we show that the branching factor of every path in $\mathcal{T}$ can be at most $r+1$. Suppose, for the sake of contradiction, there exists a path $\sigma\in\{-1,1\}^{T}$ whose branching factor is $k>r+1$. We are guaranteed an increasing sequence of time points $t_{1},t_{2},\ldots,t_{k}$ in $\{0,\ldots,T-1\}$ such that $\mathcal{T}_{t_{i}+1}((\sigma_{\leq t_{i}},-1))\neq\mathcal{T}_{t_{i}+1}((\sigma_{\leq t_{i}},+1))$ for all $i\in[k]$. Since $t_{2}\geq 1$, by our argument above, we are guaranteed that $\mathcal{T}_{t_{i}}(\sigma_{\leq t_{i}})\notin\text{span}(\{\mathcal{T}_{0},\ldots,\mathcal{T}_{t_{i}-1}(\sigma_{\leq t_{i}-1})\})$ for all $i\in\{2,\ldots,k\}$. This further implies that the set $S:=\{\mathcal{T}_{t_{2}}(\sigma_{\leq t_{2}}),\ldots,\mathcal{T}_{t_{k}}(\sigma_{\leq t_{k}})\}$ is a linearly independent set. Note that $|S|=k-1>r$. Let $W_{\sigma}$ be the matrix such that the associated function $f_{\sigma}\in\mathcal{F}$ shatters this path $\sigma$. By the definition of the tree, we must have that $S\subseteq\text{image}(W_{\sigma})$. Since $\text{rank}(W_{\sigma})\leq r$, the set $\text{image}(W_{\sigma})$ must be a subset of a vectorspace of dimension $\leq r$. However, this contradicts the fact that $S$ contains at least $r+1$ linearly independent vectors in $\mathbb{R}^{n}$. Thus, the branching factor of every path in $\mathcal{T}$ is at most $r+1$ and $\operatorname{B}(\mathcal{T})\leq r+1$. Since $\mathcal{T}$ is arbitrary, we have $\operatorname{C}_{T}(\mathcal{F})\leq r+1$.
Proof E.2.
@@
(of (ii) in Theorem 3.6) We claim that it is without loss of generality to consider the function class $\mathcal{G}=\{x\mapsto Mx\,(\text{mod }2)\,:\,M\in\{0,1\}^{n\times n}\}\subset\mathcal{F}$. Indeed, for every $f\in\mathcal{F}$, there exists a $g\in\mathcal{G}$ such that $f(x)=g(x)$ for all $x\in\mathcal{X}$. To see this, let $W\in\mathbb{Z}^{n\times n}$ be such that $f(x)=Wx(\text{mod }2)$ and fix some $x\in\{0,1\}^{n}$. Then, observe that $f(x)=Wx(\text{mod }2)=(W(\text{mod }2)x)(\text{mod }2)=Mx(\text{mod }2)$ where $M=W(\text{mod }2)\in\{0,1\}^{n\times n}$. Thus the function $g\in\mathcal{G}$ parameterized by $M$ matches $f$ everywhere on $\mathcal{X}$. We now prove that $\operatorname{C}_{T}(\mathcal{G})=n$. In the lower bound, we will use the fact that for a system of $n$ linearly independent equations with $n$ free variables defined over the integers modulo $2$, there exists a unique solution.
@@
Starting with the upper bound, fix $T\in\mathbbm{N}$ and suppose that $T\geq n$ (as otherwise $\operatorname{C}_{T}(\mathcal{G})\leq T\leq n$). Consider a trajectory tree $\mathcal{T}$ of depth $T$ shattered by $\mathcal{G}$. Let $\sigma\in\{-1,1\}^{T}$ denote an arbitrary path down $\mathcal{T}$. We first claim that for $t\geq 1$ if $\mathcal{T}_{t}(\sigma_{\leq t})$ is linearly dependent modulo $2$ on the preceding sequence of states $\mathcal{T}_{0},\mathcal{T}_{1}(\sigma_{\leq 1}),...,\mathcal{T}_{t-1}(\sigma_{<t})$, then $\mathcal{T}_{t+1}((\sigma_{\leq t},-1))=\mathcal{T}_{t+1}((\sigma_{\leq t},+1)).$ To see this, suppose that $\mathcal{T}_{t}(\sigma_{\leq t})$ is linearly dependent modulo $2$ on the preceding sequence of states $\mathcal{T}_{0},\mathcal{T}_{1}(\sigma_{\leq 1}),...,\mathcal{T}_{t-1}(\sigma_{<t})$. Then, there exists constants $a_{0},...,a_{t-1}\in\{0,1\}$ such that $a_{0}\mathcal{T}_{0}+...+a_{t-1}\mathcal{T}_{t-1}(\sigma_{<t})(\text{mod }2)=\mathcal{T}_{t}(\sigma_{\leq t}).$ Thus, every function $g\in\mathcal{G}$ consistent with the sequence $\mathcal{T}_{0},\mathcal{T}_{1}(\sigma_{\leq 1}),...,\mathcal{T}_{t-1}(\sigma_{<t})$ outputs $a_{0}\mathcal{T}_{1}(\sigma_{\leq 1})+...+a_{t-1}\mathcal{T}_{t}(\sigma_{\leq t})(\text{mod }2)$ on $\mathcal{T}_{t}(\sigma_{\leq t})$, implying that the paths $(\sigma_{\leq t},-1)$ and $(\sigma_{\leq t},+1)$ are shattered only if $\mathcal{T}_{t+1}((\sigma_{\leq t},-1))=\mathcal{T}_{t+1}((\sigma_{\leq t},+1))=a_{0}\mathcal{T}_{1}(\sigma_{\leq 1})+...+a_{t-1}\mathcal{T}_{t}(\sigma_{\leq t})(\text{mod }2).$ As a consequence, for $t\geq 1$, if $\mathcal{T}_{t+1}((\sigma_{\leq t},-1))\neq\mathcal{T}_{t+1}((\sigma_{\leq t},+1))$, then $\mathcal{T}_{t}(\sigma_{\leq t})$ is linearly independent modulo $2$ of its preceding states. Next, we claim that there can be at most $n-1$ timepoints $t_{1},...,t_{n-1}\in[T-1]$ such that $\mathcal{T}_{t_{i}}(\sigma_{\leq t_{i}})$ is linearly independent of its preceeding sequence of states $\mathcal{T}_{0},\mathcal{T}_{1}(\sigma_{\leq 1}),...,\mathcal{T}_{t_{i}-1}(\sigma_{<t_{i}})$. Indeed, suppose for sake of contradiction there exists $n$ timepoints $t_{1},...,t_{n}\in[T-1]$ such that $\mathcal{T}_{t_{i}}(\sigma_{\leq t_{i}})$ is linearly independent of its preceeding states. Then, note that the following set of $n+1$ states $\{\mathcal{T}_{0},\mathcal{T}_{t_{1}}(\sigma_{\leq t_{1}}),...,\mathcal{T}_{t_{n}}(\sigma_{\leq t_{n}})\}$ are linearly independent modulo $2$. This is a contradiction since $\mathcal{X}$ is $n$-dimensional. Combining the two claims, we get that
@@
Since $\sigma\in\{-1,1\}^{T}$ is arbitrary, we get that $\operatorname{B}(\mathcal{T})\leq n$. Finally, since $\mathcal{T}$ is arbitrary, we have that $\operatorname{C}_{T}(\mathcal{F})\leq n$ which completes the proof of the upper bound.
To prove the lower bound, let $\mathcal{A}$ be any deterministic algorithm. Let =$\mathcal{E}=\{e_{0},...,e_{n-1}\}\subset\mathcal{X}$ be the standard basis over $\mathbb{R}^{n}$. Consider the stream where we pick $x_{0}=e_{0}$, for all $t\in[n-2]$, we pick $x_{t}\in\mathcal{E}\setminus\{x_{0},...,x_{t-1},\mathcal{A}(x_{<t})\}$. Let $e=\mathcal{E}\setminus\{x_{0},...,x_{n-2}\}$ be the remaining basis. Pick $x_{n-1}\in\{e,e+e_{0}\}\setminus\{\mathcal{A}(x_{<n-1})\}$. Finally, pick $x_{n}\neq\mathcal{A}(x_{<n})$. We can define a stream using $\mathcal{A}$ because it is deterministic and thus can be simulated before the game begins. Next, we show that $x_{0},...,x_{n}$ is a realizable stream. This follows from the fact that $x_{0},...,x_{n-1}$ are linearly independent by definition, and thus there exists a function $g\in\mathcal{G}$ such that $g(x_{t-1})=x_{t}$ for all $t\in[n]$. Moreover, by definition of the stream, $\mathcal{A}$ makes a mistake in every round. Thus, $\operatorname{M}_{\mathcal{A}}(T,\mathcal{F})\geq n$. Using the fact that $\operatorname{M}_{\mathcal{A}}(T,\mathcal{F})\leq\operatorname{C}_{T}(\mathcal{F})$ completes the proof.
Proof E.3.
(of (iii) in Theorem 3.6) Starting with the lower bound, it suffices to show that
$\inf_{\text{Deterministic }\mathcal{A}}\operatorname{M}_{{\mathcal{A}}}(T,\mathcal{F})\geq n.$
@@
So, we obtain $\mathbbm{1}\{W^{\star}x_{t-1}>0\}=x_{0}-(e_{i_{1}}+\ldots+e_{i_{t}})>0=x_{t}$. Finally, since $x_{n-1}=e_{i_{n}}$ and $W^{\star}e_{i_{n}}=x_{n}$, we have $\mathbbm{1}\{W^{\star}x_{n-1}>0\}=x_{n}$. By the definition of the stream, $\mathcal{A}$ makes mistakes in every round on this stream. Thus, $\operatorname{M}_{{\mathcal{A}}}(T,\mathcal{F})\geq n$.
As for the upper bound, since |$|\{0,1\}^{n\times n}|=2^{n^{2}}$, we have $\operatorname{L}(\mathcal{F})\leq\log(2^{n^{2}})=n^{2}$. Thus, by Theorem 3.5 (iii), we must have $\operatorname{C}_{T}(\mathcal{F})\leq n^{2}$.
Appendix F Proof of Theorem 4.1
Proof F.1.
@@
where expectation is taken with respect to the randomness of $\mathcal{B}$ and $R$ is the expected regret of the $\mathcal{B}$. This shows that $\operatorname{MR}_{\mathcal{A}}(T,\mathcal{F})\leq R$. Taking $\mathcal{B}$ to be $\mathbb{A}_{\text{AG}}$ defined in (Hanneke et al., 2023, Section 3), Theorem 4 in (Hanneke et al., 2023) implies that $R\leq\operatorname{L}(\mathcal{F})+\sqrt{T\,\operatorname{L}(\mathcal{F})\log T}$, matching the claimed upper bound in Theorem 4.1.
Proof F.2.
(of lower bound of $\frac{\operatorname{L}(\mathcal{F})}{18}$ in Theorem 4.1) Define $d:=\operatorname{L}(\mathcal{F})$ and let $\mathcal{T}$ be the Littlestone tree of depth $d$ shattered by $\mathcal{F}$. Let $(\mathcal{T}_{0},\ldots,\mathcal{T}_{d-1})$ be the sequence of node-labeling functions and $(Y_{1},\ldots,Y_{d})$ be the sequence of edge labeling functions of the shattered tree $\mathcal{T}$. To construct a stream, take $r=\lfloor\frac{d+2}{3}\rfloor$ and define a random sequence $\{n_{i}\}_{i=1}^{r}$ such that
@@
$\operatorname{MR}_{\mathcal{A}}(T,\mathcal{F})\geq\frac{7r-4}{6}-(r-1)=\frac{7r-4-6r+6}{6}=\frac{r+2}{6}\geq\frac{(d+2)/3-1+2}{6}\geq\frac{d}{18}.$
This completes our proof.
Proof F.3.
@@
$\begin{split}\operatorname{MR}_{\mathcal{A}}(T,\mathcal{F})&=\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{A}(x_{<t})\neq x_{t}\}\right]-\mathbb{E}\left[\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\mathbbm{1}\{f(x_{t-1})\neq x_{t}\}\right]\geq\frac{1}{2}\mathbb{E}\left[\left|\sum_{t=1}^{T}\sigma_{t}\mathbbm{1}\{x_{t-1}\notin S\}\right|\right].\end{split}$
We now lower bound the Rademacher sum above by closely following the proof of Khinchine’s inequality (Cesa-Bianchi and Lugosi, 2006, Lemma 8.1).